Please help.
Find b^2 _4ac and the number of real solutions to each equation.
1. 9m^2 + 16 = 24m
2. -3x^2 + 7x =0

Answer by edjones(8007)

ax^2+bx+c=0
.
9m^2 + 16 = 24m
9m^2-24m+16=0 Now it is in the correct form.
b^2-4ac
=24^2-(4*9*16)
=576-576
=0 There is one real answer because b^2-4ac is zero.
.
-3x^2+7x=0 There is no constant, c so c=0.
b^2-4ac
=7^2-(4*-3*0)
=49-0
=49 There are 2 real solutions because b^2-4ac is positive.
.
Ed